# biblio.bib

### v1

@@ 3.2 Inverse Participation Ratio
We thus turn on investigating the IPR for graphs highly connected on a global scale, after tuning the tendency to clusterize and the rewiring probability. In so doing, we aim to check whether a parameter’s range exists, for which the memory-preserving property remains protected for a significant space and temporal window. Therefore, we now revert back to simulations performed over longer-enough temporal windows, so that information or energy has spread throughout the network, to evaluate the participation of the basis vectors to the initial state’s temporal evolution (given by its inverse). To explore the whole physics landscape, we now simulate a set of 10 graphs with size $L=33$, for each pair of generative parameters $(k,\beta)$ with $k=1,2,\dots,15$ (maximum possible $k=16$, equivalent to the all-to-all situation) and $\beta=0.1,0.2,\dots,1$. For each case study, we take the average value of the IPR in a temporal window chosen such that the excitation could have the time to travel across the whole graph: specifically, we simulate the temporal evolution for the period $tJ/2\in[0,100]$ and compute the average for times in the interval $tJ/2\in[90,100]$. Then, for each given pair $(k,\beta)$ we compute the average of the provided time averages for each graph.
@@
We summarize the results in Fig. 6. The boxplot displays the average IPR as a function of the number of initial neighbors $2k$, for different values of the rewiring probability $\beta$. We see that localization, i.e. higher IPR values, is favored for extremal values of $k$ (1 and 15). The lower extreme corresponds to the obvious case of low connectivity, specifically similar to a nearest-neighbor connection. The latter, instead, corresponds to the case of highly connected graphs with significant interference effects. Delocalization manifests instead at intermediate connectivity conditions. The rewiring probability $\beta$ does not significantly influence the average value, essentially dominated by the number of neighbors. We associate variations in this behavior with the limited number of sampled graphs.
@@
We now represent in Fig. 6$(b)$ the same IPR results in a scatter plot where the graphs are characterized by their average distance and average clustering (see Appendix) instead. We see that relatively localized long-time states are possible away from the limit of all-to-all connected graphs, that is with average clustering and average distance equal to 1. Quite interestingly, non-trivial steady states emerge not only in the highly connected cases but also in sparsely connected graphs with high-average distance, due to the slow propagation of excitation and small interference. This behavior heralds the potential for a memory-preserving property of the initial quench even away from the all-to-all case, should a certain decoding protocol be provided. In contrast, with $k$ values large compared to an all-to-all case with $L(L-1)/2$ (see orange arrow), signatures of cooperative shielding appear, typical of complete connection. The difference between the finite IPR values (green to light-blue symbols) shown in the top-left and in the bottom-right part of Fig. 6$(b)$ is better envisioned by inspecting Fig. 6$(c)$. Here we show a typical IPR vs. time $tJ/2$ behavior (blue line) for parameter values in the bottom-right part of $(b)$. The IPR signal oscillates around an average that is clearly above the threshold $1/L$ of complete delocalization, though the average value is lower and irregular with respect to the IPR in a corresponding all-to-all case (yellow). The cascade model prediction (red) only provides a good fit at short times. This is due to the fact that the cascade model neglects all backflow processes. Thus, we might infer that the emerging localization in this region is enhanced by quantum interference effects among different paths, in a mechanism similar to Anderson localization [alet2018_mbl].
@@
Figure 6: Effect of network topology on the long-time average IPR. For each pair $(k,\beta)$ of initial conditions in the graph generation, the computed average IPR in the time window $tJ/2\in[90,100]$, to ensure that the excitation could travel across the whole graph. $(a)$ Long-time averaged IPR vs. the number of initial neighbors $2k$ normalized to $L-1$. Each $2k/(L-1)$ value represents the median (horizontal bar inside the boxes) or quartiles and outliers (crosses). Localization is seen to occur at extremal values of $k$, i.e. nearest-neighbor and all-to-all type of connectivity. Parameter values: $k=1,\dots,15$, $\beta=0.1,0.2,\dots,1$, $L=33$, 20 samples for each couple $(k,\beta)$. $(b)$ Long-time averaged IPR in the form of a scatter plot in the parameter space of average clustering $\mathcal{C}$ and average distance $\mathcal{D}$. The arrow indicates increasing values of $k$. Notice that, away from all-to-all connectivity with $\mathcal{C}=\mathcal{D}=1$, some degree of localization is still possible for more regular networks, i.e. sparse graphs with large $\mathcal{D}$ and intermediate to very low values of clustering $\mathcal{C}$. We see this in panel $(c)$ showing a typical graph of IPR vs. dimensionless time $tJ/2$ (blue line) for parameter values $k=1,\beta=0.5,\mathcal{D}=4.8,\mathcal{C}=0$ in the bottom-right part of $(b)$, corresponding to the network in $(d)$ initialized at site $18$, as depicted. The IPR displays irregular oscillations around an average (dot-dashed line) that is well above the threshold $1/L$ of complete delocalization (dashed line). For comparison, the IPR signal is shown for the all-to-all case (yellow) and from the cascade model (red).
@@ 4 Discussion and conclusions
This work is motivated by the relevance of complex quantum networks in the modeling of biological systems and as powerful tools to extract from many-body systems relevant information connected to emerging entanglement properties or topological effects. We have therefore addressed the transport properties of quantum networks as governed by their connectivity graph and topology and by the range of the couplings connecting the nodes, properties that can be implemented in current quantum technology platforms [yagomalo2024review]. We have focused on networks regulated by the XXZ Hamiltonian, that is a paradigmatic model with a variety of applications [franchini2017integrable]. Motivated by the findings in [YagoMalo2023], we have also focused on the property of spatial localization of initial quenches introduced in the network, and the robust preservation of their memory over long time, even in the out-of-equilibrium quantum network, irrespective of where and when the excitations are switched on, and of their number. This peculiar phenomenon essentially results into a spatiotopic representation of the number of frequencies in the power spectrum of the network, that are as many as the number of excitations initially injected. Conversely, the power spectrum maps the space and time location of the initial excitations. Even more interestingly, the uncertainty in the number of excitations that the quantum network can map, linearly increases with the number itself. This behavior is shown in [YagoMalo2023] to elegantly account for the so-called Weber’s law for numerosity perception in neuroscience problems, after mapping the essential functions of the classical neuronal network into an open quantum network (resulting into the XXZ model), and in contrast with the poor performance of conventional biologically inspired models like artificial neural networks. In fact, this memory preserving property can be of much interest for concrete implementations with current quantum technologies.
@@
To gain insights on this complex matter, we have adopted a methodology that progressively develops understanding from the simpler to harder cases: we start from the all-to-all connectivity with constant coupling among the nodes, then introducing power-law interactions in the regular network, and eventually passing to irregular, complex networks. The methodology employes combined analytical and numerical methods, revolving around the cooperative shielding concept and quantum cascade model [santos2016cooperative] accompanied by exact diagonalization.
@@
Three main results emerge from our analysis. First, the memory preserving properties of the all-to-all connected network have been traced back to the symmetry structure of the Hamiltonian via the concept of cooperative shielding. Both the one-to-one correspondence of the number of excitations with the number of frequencies in the power spectrum and the number uncertainty have been analytically explained, the former as a consequence of the Wigner-Eckart theorem to get the allowed quantum numbers from the usual angular momentum composition rules. In any event, employing different quantifiers in conjunction with IPR would certainly be a best suited strategy for a thorough characterization of quantum information transport, including e.g. entanglement entropy or out-of-time ordered correlation functions.
Second, standing on the insight gained with all-to-all connectivity, but now predicting the early-time behavior with the cascade model, a similar understanding has been developed for irregular networks, by varying either the links strength according to a power law, or the network architecture via small world networks. Distinct IPR time-evolution behaviors can be observed as in Fig. 4(d), from complete localization to complete delocalization, also within individual nodes of a single complex network. The evolution between the two extremes deserves further systematic analytical and numerical investigation, to characterize its nature and explore its predictability.
Finally, our analysis leads to characterize the long-time behaviour of irregular networks to study signals of memory preservation of excitations, that is excitations that to some degree do not delocalize nor thermally relax. Quite interestingly indeed, this memory-preserving property emerges to be better protected not only in densily connected graphs close to all-to-all conditions but also in sparse and more irregular ones, i.e. with increasingly larger average distances and intermediate-to-low clustering conditions. While the average IPR value is lower and the time signal irregular with respect to the corresponding all-to-all case, effects from the initial quench are stronger in such regimes, likely due to smaller quantum interference effects. Overall, this memory-preserving property is seen to break already at short times, after introducing distance-dependent irregularities, either in the form of power-law changing links in a regular network or of large average distances among the nodes. However, we find parameter windows providing effective local regularity for the degeneracy to be protected at longer times and cooperative shielding be at work.
This work opens a number of interesting future directions. First, the case of more excitations initially injected is to be analyzed, though no significant surprises are to be expected  [YagoMalo2023, Civolani2024].
Second, it is known that the Weber’s law in neuroscience holds as well for time and space intervals, provided they are above a certain threshold. Besides, in neuroscience it is experimentally known [Morrone:2007] that the perception of numbers and of time and space intervals are interconnected with each other, suggesting that there is only one area of the brain dedicated to such an integrated mechanism. Based on a biomimicry argument, we therefore expect that similar properties can be encoded in our quantum XXZ network, envisioning how to concretely implement them in a quantum technology.
Third, the results in Fig. 6(b) suggest that the articulated behaviour of the complex quantum network might be more simply described through a universal behavior, according to one single parameter embodying all possible connectivity measures. More extended simulations and analysis can answer this question.
@@ Acknowledgements
[…]
[heading=bibintoc]
Appendix
Inverse Participation Ratio
We provide here a compact description of the Inverse Participation Ratio (IPR), used as a quantifier of the (de)localization of an initial excitation. Given a density matrix $\hat{\mathcal{\varrho}}(t)$ and the subspace of the $L$-dimensional Pauli basis $\mathcal{P}_{1}$ describing single-excitation localized states $\{|s\rangle\}$, the IPR is then the sum of $\hat{\mathcal{\varrho}}$ trace elements squared:
@@
The IPR is always comprised between $1/L$ and 1, corresponding to the limits of complete de-localization, i.e. infinite temperature case $\hat{\mathcal{\varrho}}=\hat{\mathcal{\mathbb{I}}}/L$, and perfect localization of the excitation in a state of our eigenbasis $|s_{*}\rangle$, respectively. Put it another way, the Participation Ratio (PR), which is the IPR’s inverse, indicates how many elements of the chosen basis “significantly” participate to a pure state’s temporal evolution, the lower meaning higher localization.
Characterizing complex graphs: average distance and clustering

### v2

@@ 3.5 Infinite-time localization depending on graph properties
We thus turn on investigating the IPR for graphs highly connected on a global scale, after tuning the tendency to clusterize and the rewiring probability. In so doing, we aim to study the spreading of information and its dependence on the system parameters. Therefore, we now revert back to simulations performed over longer-enough temporal windows, so that information or energy has spread throughout the network, to evaluate the participation of the basis vectors to the initial state’s temporal evolution (given by its IPR). To explore the whole physics landscape, we now simulate a set of 20 graphs with size $L=33$, for each pair of generative parameters $(k,\beta)$ with $k=1,2,\dots,15$ (maximum possible $k=16$, equivalent to the all-to-all situation) and $\beta=0.1,0.2,\dots,1$. For each case study, we take the long time average value of the IPR, computed for an initial site chosen at random, in a temporal window chosen such that the excitation could have the time to travel across the whole graph: specifically, we simulate the temporal evolution for the period $tJ/2\in[0,100]$ and compute the average for times in the interval $tJ/2\in[90,100]$.
@@
We summarize the results in Fig. 7. The boxplot, Fig. 7(a)(a), displays the average IPR as a function of the number of initial neighbors $2k$, for different values of the rewiring probability $\beta$. We see that localization, i.e. higher IPR values, is favored for extremal values of $k$ (1 and 15). The lower extreme corresponds to the obvious case of low connectivity, specifically similar to a nearest-neighbor connection. The latter, instead, corresponds to the case of highly-connected graphs with significant interference effects. Delocalization manifests instead at intermediate connectivity conditions. The rewiring probability $\beta$ does not significantly influence the average value, essentially dominated by the number of neighbors. We associate variations in this behavior with the limited number of sampled graphs.
@@
We now represent in Fig. 7(b)(b) the same IPR results in a scatter plot where the graphs are characterized by their average distance and average clustering instead, see Appendix C. We see that relatively localized long-time states are possible away from the limit of all-to-all connected graphs, that is with average clustering and average distance equal to 1. Quite interestingly, non-trivial steady states emerge not only in the highly connected cases but also in sparsely connected graphs with high-average distance (lower-right part of the plot), due to the slow propagation of excitation and small interference. This behavior heralds the potential for a memory-preserving property of the initial quench even away from the all-to-all case, should a certain decoding protocol be provided. In contrast, with $k$ values comparable to the all-to-all case with L(L1)/2L(L-1)/2 (upper-left part), signatures of cooperative shielding appear, typical of complete connection. The difference between the finite IPR values (green to light-blue symbols) shown in the top-left and in the bottom-right part of Fig. 7(b)(b) is better envisioned by inspecting Fig. 7(c)(c). Here we show a typical IPR vs. time $tJ/2$ behavior (blue line) for parameter values in the bottom-right part of $(b)$ for $k=1$. The IPR signal oscillates around an average (dot-dashed horizonatla line) that is clearly above the threshold $1/L$ (dashed horizontal line) of complete delocalization, though the IPR has a lower value with respect to the IPR in a corresponding all-to-all case (yellow). The cascade model prediction (red) only provides a reasonable fit at all times. The differences are due to the fact that the cascade model neglects all backflow processes. Thus, we might infer that the emerging localization in this region is enhanced by quantum interference effects among different paths, in a mechanism similar to Anderson localization [80]. Finally we note that the question of whether cooperative shielding survives in more general networks deserves further study which requires a scaling analysis as a function of the system size. This topic will be analyzed in a future work.
@@
Figure 7: Effect of network topology on the long-time average IPR. For each pair $(k,\beta)$ of initial conditions in the graph generation, the computed average IPR in the time window $tJ/2\in[90,100]$, to ensure that the excitation could travel across the whole graph. Each pair is tested 20 times and each time the initial site is chosen at random. $(a)$ Long-time averaged IPR vs. the number of initial neighbors $2k$ normalized to L1L-1. The horizontal dashed line on top is the average IPR for the all-to-all case. For each 2k/(L1)2k/(L-1) value, boxes represent the data’s interquartile range, i.e. all points whose mean IPR belong between the 25th and 75th quartile, and have a horizontal line marking the median; whisker lines extend up to 1.5 times the interquartile range outside of boxes; crosses depict the outliers, or points that are not included in neither boxes or whiskers; finally, the red diamonds display the average IPR sampling. Localization is seen to occur at extremal values of $k$, i.e. nearest-neighbor and all-to-all type of connectivity. Parameter values: $k=1,\dots,15$, $\beta=0.1,0.2,\dots,1$, $L=33$. $(b)$ Long-time averaged IPR in the form of a scatter plot in the parameter space of average clustering $\mathcal{C}$ and average distance $\mathcal{D}$ (data having $\mathcal{C}=0$ is not shown). The general trend going from lower-right to upper-left denote increasing values of $k$. Notice that, away from all-to-all connectivity with $\mathcal{C}=\mathcal{D}=1$, some degree of localization is still possible for more regular networks, i.e. sparse graphs with large $\mathcal{D}$ and intermediate to very low values of clustering $\mathcal{C}$. We see this in panel $(c)$ showing a typical graph of IPR vs. dimensionless time $tJ/2$ (blue line) for parameter values $k=1,\beta=0.5,\mathcal{D}=7.0,\mathcal{C}=0$ in the bottom-right part of $(b)$, corresponding to the network in $(d)$ initialized at site $17$, as depicted. The IPR displays irregular oscillations around an average (dashed line) that is well above the threshold $1/L$ of complete delocalization. For comparison, the IPR signal is shown with the respective cascade model (red).
3.6 Power-law-dependent hopping
A different way to break space regularity in networks is to consider exchange coupling that vary with physical distance, specifically characterized by a power law, i.e. $J_{ij}\propto\text{dist}(i,j)^{-\alpha}$ in Eq. (1) which in the case of PBC is equivalent to min(|ij|,L|ij|)(|i-j|,L-|i-j|). This test case constitutes a good bench to showcase cooperative shielding, introduced in Section 1, when considering different initial conditions: in fact, on the one hand increasing values of the power law coefficient cause shorter ranges for the interactions (as was for diverse combinations of $k$ and $\beta$ for small-worlds), and on the other hand each site is equivalent to the others (unlike the previous varying network case). Here we analyze the memory effects in such models and the dynamics focusing on the cooperative shielding effects.
3.6.1 Memory effects
Here we analyze the memory-preserving effect discussed in Section 2.1 in networks with power-law interactions, starting with a single excitation and taking the Fourier transform of $S^{z}_{\text{initial}}(t)$, see Fig. 8(a)(a). As we increase $\alpha$ the high-energy frequencies are reduced but they still persist. Also in this case, this result could allow us to extend the notion of memory-preserving effect to generic range of hopping.
3.6.2 Dynamics/shielding
[…]
$\Delta\tau_{1/2}=\tau_{1/2}^{a.q.s.}-\tau_{1/2}^{s.s.}\,.$ (15)
The results are displayed in Fig. (9)$(b)$, where we obtain $\Delta\tau_{1/2}$ for different values of $\alpha$ and different scales or number of sites $L$. We first notice that the stopping time difference is a decreasing function of $\alpha$ and reaches the lower bound corresponding to the nearest-neighbor network case already for $\alpha\gtrsim 5$.
Figure 8: Memorization in all-to-all connected graphs with varying interaction range. $(a)$ Average local $z$-magnetization at the initial site for different power-law exponents $\alpha$ regulating the range of the couplings. The network possesses 33 sites and J/2J/2=1. (b)(c)(b)-(c) Corresponding power spectra for the signals depicted in Panel $(a)$ with $\alpha$=0.1,0.5; for comparison, the case $\alpha=0$ is shown in Fig. 4(b)(b). The data has been detrended before the Fourier transform by removing the zero-frequency component.
The results shown in this section are consistent with what was claimed in [71] where it was shown that single excitation transport in presence of long range hopping is highly dependent on the initial conditions. The results presented here confirm that this is a general effect of long-range hopping models. Indeed for short-range hopping, $\alpha\gg 1$ no dependence on the initial condition is shown, see Fig. 9(a)(a). Note that the antisymmetric initial state has a large overlap with the shielded subspace [71], even if it does not lie completely in the shielding subspace and this could explain the lack of cooperativity (dependence on the size of the system) shown in Fig. 9(b)(b). The data has been detrended before the Fourier transform by removing the zero-frequency component.
[…]
@@ 4 Discussion and conclusions
This work is motivated by the relevance of complex quantum networks in the modeling of biological systems and as a powerful tool to extract relevant information for many-body systems. We have therefore addressed the transport properties of quantum networks as governed by different parameters such as the connectivity graph, the topology and the range of the couplings connecting the nodes. These parameters can be tuned in current quantum technology platforms [14]. We have focused on networks described by the XXZ Hamiltonian, that is a paradigmatic model with a variety of applications [74]. To gain insights on this complex matter, we have adopted a methodology that progressively develops understanding from the simpler to harder cases: we start from the all-to-all connectivity with constant coupling among the nodes, then introducing power-law interactions in the regular network, and eventually passing to irregular, complex networks. The methodology employs combined analytical and numerical methods, revolving around the cooperative shielding concept and quantum cascade model [70] accompanied by exact diagonalization. The main focus of this manuscript is on the effect of long range interactions and connectivity on the dynamics of the network. We show that long range coupling in highly connected networks imposes strong constraints on the dynamics which manifest themselves in the memory-preserving and cooperative shielding effect [70, 71].
Motivated by the findings in [5], we have analyzed the robust preservation of the memory of the number of initial excitations over long times. This peculiar phenomenon essentially results into a spatiotopic representation of the number of frequencies in the power spectrum of the network, that are as many as the number of excitations initially injected. Even more interestingly, the uncertainty in the number of excitations that the quantum network can map, linearly increases with the number itself. This behavior is shown in [5] to elegantly account for the so-called Weber’s law for numerosity perception in neuroscience problems, after mapping the essential functions of the classical neuronal network into an open quantum network (resulting into the XXZ model), and in contrast with the poor performance of conventional biologically inspired models like artificial neural networks. In fact, this memory preserving property can be of much interest for concrete implementations with current quantum technologies.
@@
Three main results emerge from our analysis. First, the memory preserving properties of the all-to-all connected network have been traced back to the symmetry structure of the Hamiltonian. Both the one-to-one correspondence of the number of excitations with the number of frequencies in the power spectrum and the number uncertainty have been analytically explained, the former as a consequence of the Wigner-Eckart theorem to get the allowed quantum numbers from the usual angular momentum composition rules. The memory preserving effect has been analyzed also in more irregular networks, such as small world network and power law connected network. We have shown that signatures of the memory preserving behavior also persists in more irregular networks. This finding opens the possibility to generalize this effect in a wider range of network topology.
Together with the memory-preserving effect we have also analyzed the spreading of a single excitations in different networks. We have shown that due to cooperative shielding, counter-intuitively, the more connected is the network or the more long ranged are the couplings, the less the initial excitation spreads through the network. Clearly, a lower excitation spreading can be obtained also in very low connected network, see Fig. (7). These findings points out to a possible optimal connectivity for the excitation mobility, or even the possible existence of a sort of percolation transition driven by the connectivity. This topic will be explored in a future work.
Finally, new evidence of the cooperative shielding effects have been shown in power-law connected networks, where due to cooperative shielding a strong dependence on the initial conditions and on the range of the interaction have been shown, see Fig (9).
The results presented in this manuscript open different new paths in the study of highly connected networks.
@@ Acknowledgements
Appendix A Degeneracy formula for the XX, all-to-all connected energy levels
In Section 2 we introduced the recursive formula $f(l,L)$ for computing the degeneracy of a certain value of the $\hat{\mathcal{M}}^{2}$ quantum number $l$ for a given system size $L$; here we prove the result. As the Hamiltonian in Eq. 2 is written in terms of global spin operators, the Hilbert space of possible spin configurations (Pauli strings) can be decomposed into irreducible representations of spin $l$, each one with dimension $2l+1$: this means that the total Hilbert space is a direct sum of such representations.
Therefore, the question is to calculate the multiplicity of the spin-$l$ representation, given by $f(l,L)$. We build our understanding by induction starting from a single spin-1/2 (the base case): there is only one possible value for $l$, namely 1/2, thus $f(1/2,1)=1$ and any other value of $l$ gives 0. Suppose now that we have system size L1L-1 and we add a new spin-1/2: the theory of composition of angular momenta tells us that each original spin-$l^{\prime}$ from the L1L-1 system will give result to two new subspaces, one with spin-$l=l^{\prime}+1/2$ and the other with spin-$l=l^{\prime}-1/2$ (the latter only if $l^{\prime}\neq 0$). In other words, $f(l,L)$ is the sum of multiplicities of spin-l1/2l-1/2 and spin-$l+1/2$ representations at the step L1L-1, as expressed by the recursive formula f(l,L)=f(l1/2,L1)+f(l+1/2,L1)f(l,L)=f(l-1/2,L-1)+f(l+1/2,L-1). A visual construction is given in Fig. 10 for system sizes from 1 to 4: at each row we count how many times the allowed value for $l$ appears, which accounts for the spin-$l$ multiplicity. For example, in the case $L=4$, the allowed value for $l$ are 0, 1, 2, so that the respective multiplicities $f(l,L)$ are 2, 3, 1. Once the spin-$l$ multiplicity known, the actual level degeneracy, dependent on $l$ and $|m|$, can be computed as done in the text.
Figure 10: Consecutive composition of spin-1/2 angular momenta. The rows of the triangle represent increasing system sizes from 1 to 4. The nodes are the possible values of $l$, obtained either by adding (blue, continuous arrows) or subtracting (red, dashed arrows) 1/2 from the element above.
Appendix B Inverse Participation Ratio
We provide here a compact description of the Inverse Participation Ratio (IPR), used as a quantifier of the (de)localization of an initial excitation. Given a density matrix $\hat{\mathcal{\varrho}}(t)$ and the subspace of the $L$-dimensional computational basis $\mathcal{P}_{1}$ describing single-excitation localized states $\{|s\rangle\}$, the IPR is then the sum of $\hat{\mathcal{\varrho}}$ trace elements squared:
@@ Appendix B Inverse Participation Ratio
The IPR is always comprised between $1/L$ and 1, corresponding to the limits of complete delocalization, i.e. infinite temperature case $\hat{\mathcal{\varrho}}=\hat{\mathcal{\mathbb{I}}}/L$, and perfect localization of the excitation in a state of our eigenbasis $|s_{*}\rangle$, respectively. Put it another way, the Participation Ratio (PR), which is the IPR’s inverse, indicates how many elements of the chosen basis “significantly” participate to a pure state’s temporal evolution, the lower meaning higher localization.
Appendix C Characterizing complex graphs: average distance and clustering